OpenStudy (anonymous):

Jules kicks a soccer ball off the ground and into the air with an initial velocity of 25 feet per second. Assume the starting height of the ball is 0 feet. Approximately, how long does it take until the soccer ball hits the ground again? 0.6 sec 0.8 sec 1.6 secs 2.8 secs

OpenStudy (anonymous):

who has gamail

OpenStudy (anonymous):

if the question means 25 velocity vertically...then you do this... vf = vi + at 0 = 25 + (-9.8) t t = 25 / 9.8 t = 2.55 that's the time it takes to reach the maximum height multiply that by two to get the time it takes to fall back to the ground so around 5 seconds